What volume of 0.40 M NH3 solution must be added to 1Lof 0.1 m NH4Cl solution to give a buffer having pH of 10?For NH3, KB=1.8^10-5.

Grade:11

1 Answers

vikas askiitian expert
13 years ago

[H+] = 10-10

[OH-] = Kw/H+ = 10-4

moles of NH4Cl = 0.1*1=0.1

for V Litre of NH3 ,moles of NH3 = 0.4*V 

now ,

          Kb = [salt][OH-]/[base]

      1.8*10-5 = (moles of salt )*10-4/(moles of base)

    moles of base(NH3) =  10/1.8*4=1.38L